a jua kali artisan produces specially designed leather belts. the demand curve for the belts is a follows
price ksh quantity of belts
1500 200
2000 100
calculate the arc elasticity of demand and interpret
1
Expert's answer
2019-03-21T12:31:26-0400
"Kd=(Q2-Q1)\/((Q1+Q2)\/2):(P2-P1)\/((P1+P2)\/2)"

So,

Kd=(100-200)/150:(2000-1500)/1750=-2.33

This means that if price increases by 1% the demand decreases by 2.33%